GreenHab Report – February 26th

Green Hab Report  – Sol 14
Report written by: Victoria DA-POIAN (Crew Biologist)

Date : 02/26/2017

Functionality: The heater in the green hab is working well. Today was a very sunny but cold day ! Arthur checked the temperature in the GreenHab this morning. It was around 41 Celsius degrees around 11:00 AM while the temperature in the tunnel was around 6.6 Celsius degrees. He switched on the cooler on stage 3. The GreenHab temperature was 21 Celsius degrees at 6:00PM while it was 8.9 Celsius degrees in the tunnel. we watered twice the plants today.

Status: The existing seedlings in the greenhab are continuing to grow well. There are spinach, lettuce, radish, and beans growing very well in the small pots.

The lettuces are growing well too. The new lettuces I planted are growing very well too.

The Vegidair has been installed a week ago and is functionning very well. We were able to see nice sprouts of lettuce today in the Vegidair and some smaller in the similar pots I put in the GreenHab.


Commander Report – February 26th

Dear Earth,

Yesterday’s misadventures made us realize that we needed to take a
break. We had been doing our thirty-minutes physical training every
single Sol since our arrival, at 7 AM before breakfast, and almost an
EVA every morning. It was maybe too much, so today we decided to wake
up later: on Sol 14, no early workout, some sort of brunch, and an EVA
planned for the afternoon instead of the morning. That is what a
Sunday on Mars should look like.

The EVA was dedicated to emergency procedures. After a quick briefing
on lower deck, Xavier led his four EVA buddies outdoors. Mouâdh, Louis
Maller and I tested a set of first-aid positions on Xavier who was
supposed to be a fainted or injured astronaut, while Simon was taking
pictures of the scene in order to make a tutorial. The most difficult
part was to put the unconscious body of Xavier on the rear of the
rover, but after some unsuccessful tries using ropes, we managed to
put him in recovery position and carry him around the Hab’s parking

Later, we did a general cleaning of the Hab, for it to be ready when
our guest astronaut from the BBC arrives tomorrow. Indeed, we have the
privilege to have the british physicist Brian Cox joining our crew for
one Sol. We are eager to meet him!

Ad Astra!
Arthur Lillo
Commander of the emergency-ready Crew 175

Crew Photos – February 25th

Orion Nebula


Inflating the balloon


Crew walk


Balloon take off


Rovers in grey Moon


Crew in the sun


Yellow and grey moon

Journalist Report – February 25th

Journalist report, 02/25/17 – Sol 13:  Water games.

We started this sol with very bad news: the water pump we use to refill the indoor tank from the outdoor one was not responding, and we only had 4 gallons (15 L) left. Enough to flush twice, to take two showers, to drink for a bit more than one day… But we had to wash dishes, and cook, which also requires a huge amount of water because of the food we have to rehydrate. We didn’t hear any noise coming from the pump so that we believed that it was an electrical issue, or a problem coming from it, but kept in mind that the temperature during the night would have allowed it to freeze easily. We decided to first go on EVA, so that we might see if there was any problem outside.

The EVA was supposed to start with an outdoor water tank refill, but we found out that the portable pump we use for it wasn’t functioning too. It had stayed in the engineering airlock so that it might be frozen too. We aborted the refill and the EVA started by giving another try with the solar balloon. This time it went right: Simon, in charge of the experiment, leading this EVA deployed it along with Arthur, Louis and Mouâdh. They used the spot we determined before, and let it fly there for the whole EVA, while they left for “the Moon” again, this time with Arthur and his sextant, in order to let him measure coordinates in this new location. At the same time, Xavier was trying to find out what was wrong with the pump, and one hour later, it started again as usual. It must have frozen during this night, even if it was in the airlock. We had solved one problem, but the main pump was not working yet, even if we had tried to put hot water on the outside pipe coming out of the tank.

The EVA went fine on the North, with some new beautiful shots from Louis, while the balloon was doing fine, oscillating slowly from the ground to 50 meters high. When they came back, we had prepared every pan and bowl we had in the hab, for an emergency refill. The 4 gallons we had didn’t allow us to cook the meal. The EVA members, back from their mission, spent 15 minutes and lost a lot of water filling up all the stuff we had let in the engineering airlock. About 5 minutes after all this trouble, the main pump decided to restart working… Water must have frozen in the outdoor pipe too. Anyway it ended up well even if we had to move around 10 gallons of water in the stairs (that are closer to a ladder by the way).

Tomorrow, we will begin with our last experiment: simulating emergency situations in EVA. Victoria and I will be the only one to remain in the hab in the morning, while 4 crew members will experiment it being led by Xavier.

Commander Report – February 25th

Dear Earth,

The Hab is definitely not made to host a crew for more than two weeks:
we spent the first 12 Sols of a typical two-week rotation perfectly
safe in our little haven, and now that we reached the fateful Sol 13,
it seems that the Hab is trying to kill us all… some say that it was
built on a Martian cemetery. Moreover, we feel like a small family
that would take care of an isolated hotel in the middle of the winter,
if you see what I mean.

This morning, Simon, Mouâdh, Louis Maller and I deployed the balloon
north of the Hab and then visited the grey landscapes of the Moon-like
area. Meanwhile inside the Hab, Xavier, Louis Mangin and Victoria
faced a rather critical issue: no water flowed up from the exterior
tank to the upper deck tank, and we only had 15L (4 gal) left in the
latter. In our referential, it represents 2 showers, or 2 flushes, or
a little less than what is needed to cook and wash the dishes. Water
is everything for us, particularly since most of our food needs to be
rehydrated… Fortunately, just before twelve, nature stopped playing
with our nerves and pushed the ice cubes out of the pump.
Unsurprisingly, the night was so cold that the pipes were frozen. The
melting happened just in time for the return of the brave explorers!

In the afternoon, Simon and Xavier established a set of emergency
procedures for medical issues happening outdoors. Indeed, we want to
try some first-aid moves during later EVAs, after having adapted them
to our spacesuits and vehicles.

The obscure force that seems to inhabit the Hab struck again later in
the afternoon. When one of us climbed to the loft to take a look at
the water counter, this unfortunate fellow fell violently off the
fourth bar of the ladder, because of the lack of an upper bar to hold
on to. Everyone else rushed to the injured crewmember and moved him to
the sofa. We were quick and efficient, and did not go easy on
anti-pain medication and ice. Luckily, the Fallen Astronaut is now
safe and sound. The Overlook Hote… sorry, the Hab will not trap us so

Ad Astra!

Arthur Lillo
Commander of the Shining-proof Crew 175

EVA Report – February 25th

EVA Report 25/02/2017
Crew members: Simon Bouriat (EVA leader), Louis Maller, Arthur Lillo and
Mouadh Bouayad (EVA buddies)
•       Around the Hab
•       North of the Hab (12N 518750, 4252750) if time remaining.
Vehicles used:          Deimos, 2×350 ATV
Departure time:         9:05 AM
Return time:            12:15 PM
Duration:               3 hours

One of the best EVAs I did since the beginning.
We first experimented with the solar balloon. It actually worked pretty
well despite the variable wind. We succeeded in shooting videos with the
GoPro from inside the balloon’s platform and we also put the Arduino
system to get the temperature and pressure variations. Everything went
well but because of the large ups and downs of the balloon we stayed for
almost an hour and a half to be sure that it wouldn’t break.
After that, we took Deimos and two ATVs and went to the North. We
finally reached beige moon, grey moon and then yellow moon and took
amazing pictures. Our commander tried to locate us with the sextant but
the map he printed wasn’t precise enough to get the three landmarks
needed. And unfortunately, the glasses went unexpected out of range
shortly after the beginning of the EVA. But we observed that the battery
kept them working during the entire 3 hours so we will take them again
on a next EVA.

GreenHab Report – February 25th

Green Hab Report  – Sol 13
Report written by: Victoria DA-POIAN (Crew Biologist)

Date : 02/25/2017

Functionality: The heater in the green hab is working well. Today was a very sunny day and warmer than yesterday. I checked the temperature in the GreenHab this morning. It was around 41 Celsius degrees around 11:00 AM while the temperature in the tunnel was around 3.8 Celsius degrees. I switched on the cooler (on stage 3).

The GreenHab temperature was 18 Celsius degrees at 5:00PM while it was 6 Celsius degrees in the tunnel. I then cut the cooler and watered the seedlings.

Status: The existing seedlings in the greenhab are continuing to grow well. There are spinach, lettuce, radish, and beans growing very well in the small pots. A new lettuce has been planted today. Another lettuce is not looking well, we will probably harvest it soon and plant a new one.

The Vegidair has been is functionning very well. We were able to see taller sprouts of lettuce and of « roquette » today in the Vegidair than in the similar pots I put in the GreenHab, but those in the GreenHab have also taken well.


Sol Summary – February 25th


SOL: 13
Person filling out Report: Louis MALLER, XO
Summary Title: Exploring Cold Mars
Mission Status: successful EVA in the morning, all systems go, work on experiments ongoing
Sol Activity Summary: sport, EVA, science work on different experiments
Look Ahead Plan: Tomorrow there should be an EVA aimed at testing the feasibility of safety protocols designed today
Anomalies in work: 3D printer being troubleshooted, pumps frozen in the morning (thawed later in the morning)
Weather: beautiful and sunny in the morning slight breeze, quite cold
Crew Physical Status: One crewmember fell from the ladder to the loft, bruised but nothing broken
EVA: tried to fill in the static tank at the beginning of the EVA, failed because the pump had frozen. Succesfully deployed the balloon in the plain north of the Hab. Went to Explore the Moon again, and then headed southwards towards Skyline Rim. Then we headed back to the Hab, folded up the balloon, and managed to activate the pump.

Reports to be filed:
– Commander report
– Operations report
– Journalist report
– GreenHab Report
– EVA #12 report
– EVA #13 request
– HSO report
– Astronomy Report

Commander Report – February 24th

Dear Earth,

Sol 12 was particularly cold.

This morning, Louis Mangin, Louis Maller and Simon did an EVA under
the lead of Xavier. Our Executive Officer was equipped with the
augmented reality glasses “Optinvent”, so that Mouâdh could watch on
his computer inside the Hab what Louis was seeing outdoors. However,
this helpful technology was only usable in the vicinity of the Hab
(for instance during the engineering check), because of the short
range of our WiFi connection. The rest of the EVA was dedicated to
exploration, heading to the remarkable landscapes beyond Tank Wash. An
area east of Factory Butte looked a lot like the grey hills around the
Apollo landing sites. Our four brave astronauts came back
freezing-cold and were overjoyed by the hot meal they encountered on
upper deck.

The afternoon began with card game time (the main debate here is about
what game should be played after lunch, between “la coinche”, “la
belote” and “le tarot”). Today “la coinche” won the fight, and was
followed by the usual partition of the crew in the different modules:
Victoria in the GreenHab, the two Louis and Mouâdh on the upper deck
of the Hab, Xavier, Simon and I in the Science Dome. In the meantime,
we tried to revive the 3D printer which is still reluctant to our
efforts. I am optimistic though, given the inventivity of my team.

Ad Astra!

Arthur Lillo
Commander of the cold-resistant Crew 175

Crew Photos – February 24th

12N 425300 51800 02242017 Crew on hill 2
12N 515000 4254000 02242017 ATVs on Yellow Moon
02242017 Calibrating the printer
02242017 Connected glasses connected to PC
02242017 Refilling the tank
12N 425300 51800 02242017 Crew on hill
12N 515000 4254000 02242017 From Yellow to White moon
12N 515000 4254000 02242017 Grey Moon
02242017 Deploting the balloon